Can you please solve this for me?: I have to solve this using factoring.
Originally a rectangle was twice as long as it was wide. When 4m were added to its length and 3m subtracted from its width, the resulting rectangle had an area of 600m2(squared). Find the dimensions of the new rectangle.

Answer by stanbon(75887)

Original measurements: length=2x width=x
New measurements: length=2x+4m width=x-3m
EQUATION:
(2x+4m)(x-3m)=600
2x^2-2mx-12m^2=600
x^2-mx-6m^2=300
You have an equation, but it has two variables.
Need another equation to solve for x and m
Cheers,
Stan H.